For any real number x, prove that x(x+1)(x+2)(x+3)+1 ≥ 0
1
Expert's answer
2019-05-17T05:43:59-0400


x(x+1)(x+2)(x+3)+1=(x+1)(x+2)[x(x+3)]+1=(x2+2x+x+2)[x(x+3)]+1=

(x2+3x+2)[x(x+3)]+1=[x(x+3)+2][x(x+3)]+1=

[x(x+3)]2+2[x(x+3)]+1=[x(x+3)+1]2.

So

x(x+1)(x+2)(x+3)+1 = [x(x+3)+1]2.

Consequently x(x+1)(x+2)(x+3)+1 is a perfect square.


As x(x+1)(x+2)(x+3)+1 is a perfect square, a perfect square for real number

is always non-negative ( ≥ 0).